For what values of t and s does the equality

(-(2t + 3),s+t) = (2s – 6t, 1 - 3t)

hold true?

t =  and s =

At these values, the resulting vector is
Answer #1

То елЈ6+), со +5) 5 +E) (-(20+), 5+t) = (25-67 1-3t) >) - (2x+3) = 25-6€ -Band 5+t = -3t from© - 25 - 3 = 25-64 45-25= 3 from